For a child (someone under the age of 18) to be considered “gifted” they must have a top 10% IQ. If IQ scored for someone under the age of 18 is normally distributed with a mean of 100 and standard deviation of 15, what IQ score would they need to be considered “gifted”?
Answer #1


= 100

= 15

Top 10% corresponds to area = 0.50 - 0.10 = 0.40 from mid value to Z on RHS.

Table of Area Under Standard Normal Curve gives Z = 1.28

So, we get:

Z = 1.28 = (X - 100)/15

So,

Z = 100 + (1.28 X 15)

= 119.2

So,

Answer is:

119.2